OpenStudy (anonymous):

rose had a bottle of water that was 4/5 full. during field day, she drank 2/3 of the bottle. what fraction of the bottle did she drank show your work

OpenStudy (nikato):

just multiply the fractions

OpenStudy (anonymous):

Basically you have to make the denominators the same and subtract! Then probably reduce the final fraction to it's lowest reduction.

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 } - \frac{ 2 }{ 3 } =\frac{ 4 }{ 5 }(\frac{ 3 }{ 3 }) - \frac{ 2 }{ 3 } (\frac{ 5 }{ 5 })\]

OpenStudy (anonymous):

do u understand? after that u just simplify it..

OpenStudy (nikato):

no. you multiply 2/3 of the 4/5 full bottle